# Convergence Analysis of Non-Strongly-Monotone Stochastic Quasi-Variational Inequalities

### v3

@@ Appendix C Almost Sure Convergence
$(x_{k}-\bar{x}_{k})^{T}H(x_{k})$ $\geq\tfrac{1}{\alpha_{k}}\|H(x_{k})\|^{2}+\eta\alpha_{k}(x_{k}-\bar{x}_{k})^{T}(F(x_{k})-F(\bar{x}_{k}))-\eta(F(x_{k})-F(\bar{x}_{k}))H(x_{k})$
$\geq\tfrac{1}{\alpha_{k}}\|H(x_{k})\|^{2}+\mu_{F}\eta\alpha_{k}\|x_{k}-\bar{x}_{k}\|^{2}-\tfrac{\eta^{2}L^{2}}{2}\|x_{k}-\bar{x}_{k}\|^{2}-\tfrac{\|H(x_{k})\|^{2}}{2}.$ (A7)
Where the second inequality is obtained by using the QG property, Lipschitz continuity of $F$, and implementing Young’s inequality. Next, substituting (C) and (C) into (C) yields:
$\|x_{k+1}-\bar{x}_{k}\|^{2}$ $\leq\|x_{k}-\bar{x}_{k}\|^{2}+\|H(x_{k})\|^{2}+2\alpha_{k}^{2}(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2})-2\langle x_{k}-\bar{x}_{k},H(x_{k})\rangle$
$\quad+2\langle x_{k}-\bar{x}_{k},\alpha_{k}e^{\prime}_{k}\rangle-2\langle H(x_{k}),\alpha_{k}e^{\prime}_{k}\rangle$
$\leq(1-2\mu_{F}\eta\alpha_{k}+\eta^{2}L^{2}+\alpha_{k}^{2})\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\quad+(4\alpha_{k}^{2}+1)(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2}).$
Now, by considering $\eta=\alpha_{k}$, dropping negative terms and taking the conditional expectation, one can obtain:
$\mathbb{E}[\|x_{k+1}-\bar{x}_{k}\|^{2}]$ $\leq(1+\alpha_{k}^{2}(L^{2}+1))\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\quad+2(2\alpha_{k}^{2}+1)\big{(}(\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)})^{2}+\tfrac{\alpha_{k}^{2}\nu^{2}}{N_{k}}\big{)}.$
@@
$\mathbb{E}[\|x_{k+1}-\bar{x}_{k+1}\|^{2}]$ $\leq(1+\alpha_{k}^{2}(L^{2}+1))\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\quad+2(2\alpha_{k}^{2}+1)\big{(}(\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)})^{2}+\tfrac{\alpha_{k}^{2}\nu^{2}}{N_{k}}\big{)}.$
Since $\sum_{k=0}^{\infty}\tfrac{1}{\sqrt{N_{k}}}<\infty$, by choosing $\alpha_{k}=\tfrac{1}{k^{0.5+\delta}}$ it is easy to show that the requirements of Lemma 3 are satisfied, and we can conclude that (i) $\|x_{k}-\bar{x}_{k}\|^{2}$ is a convergent sequence. Hence, the sequence $\{x_{k}\}_{k}$ is bounded, which guarantees the existence of a convergent subsequence $x_{k_{q}}\to x^{\#}$ as $q\to\infty$, for some $x^{\#}\in X$. (ii) $\sum_{k=0}^{\infty}(\tfrac{2}{\alpha_{k}}-2)\|H(x_{k})\|^{2}<\infty$ almost surely. From the definition of $H(x_{k})$, it follows that $\sum_{k=0}^{\infty}2\alpha_{k}(x_{k}-\mathcal{P}_{K(x_{k})}[x_{k}-\eta F(x_{k})])<\infty$. Since $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, we conclude that $\lim_{k\to\infty}x_{k}-\mathcal{P}_{K(x_{k})}[x_{k}-\eta F(x_{k})]\to 0$, which in turn implies that $\lim_{q\to\infty}x_{k_{q}}-\mathcal{P}_{K(x_{k_{q}})}[x_{k_{q}}-\eta F(x_{k_{q}})]\to 0$. Now considering (i) and (ii) we can show that $x^{\#}-\mathcal{P}_{K(x^{\#})}[x^{\#}-\eta F(x^{\#})]=0$, which means that $x^{\#}$ is a solution of SQVI. Therefore, we conclude that the sequence $\{x_{k}\}$ has at least one limit point, and every limit point is a solution of the SQVI almost surely. ∎

### v4

@@ Appendix C Almost Sure Convergence
$(x_{k}-\bar{x}_{k})^{T}H(x_{k})$ $=(x_{k}-\bar{x}_{k})^{T}\widehat{H}(x_{k})+(x_{k}-\bar{x}_{k})^{T}(H(x_{k})-\widehat{H}(x_{k}))$
$\geq\tfrac{1}{\alpha_{k}}\|\widehat{H}(x_{k})\|^{2}+\eta\alpha_{k}(x_{k}-\bar{x}_{k})^{T}(F(x_{k})-F(\bar{x}_{k}))$
$\qquad-\eta(F(x_{k})-F(\bar{x}_{k}))^{T}\widehat{H}(x_{k})+(x_{k}-\bar{x}_{k})^{T}(H(x_{k})-\widehat{H}(x_{k}))$
$\geq\tfrac{1}{2\alpha_{k}}\|\widehat{H}(x_{k})\|^{2}+\mu_{F}\eta\alpha_{k}\|x_{k}-\bar{x}_{k}\|^{2}-\tfrac{\eta^{2}L^{2}}{2\tau_{k}}\|x_{k}-\bar{x}_{k}\|^{2}$
$\qquad-\tfrac{\tau_{k}}{2}\|\widehat{H}(x_{k})\|^{2}-\alpha_{k}\gamma\|x_{k}-\bar{x}_{k}\|^{2},$ (A7)
where the last inequality is obtained by using the QG property, Lipschitz continuity of $F$, implementing Young’s inequality, and $|(x_{k}-\bar{x}_{k})^{T}(H(x_{k})-\widehat{H}(x_{k}))|\leq\alpha_{k}\|x_{k}-\bar{x}_{k}\|\|\mathcal{P}_{K(x_{k})}[u]-\mathcal{P}_{K(\bar{x}_{k})}[u]\|\leq\alpha_{k}\gamma\|x_{k}-\bar{x}_{k}\|^{2}$. Now, using the fact that $a^{2}\geq\frac{1}{1+\xi}b^{2}-\frac{1+1/\xi}{1+\xi}(a-b)^{2}$ we obtain $\|\widehat{H}(x_{k})\|^{2}\geq\frac{1}{1+\xi}\|H(x_{k})\|^{2}-\frac{1}{\xi}\|\widehat{H}(x_{k})-H(x_{k})\|^{2}$ which combined with (C) implies that
$(x_{k}-\bar{x}_{k})^{T}H(x_{k})$ $\geq\left(\tfrac{1}{\alpha_{k}}-\tfrac{\tau_{k}}{2}\right)\left(\tfrac{1}{1+\xi}\right)\|H(x_{k})\|^{2}-\frac{1}{\xi}\left(\tfrac{1}{\alpha_{k}}-\tfrac{\tau_{k}}{2}\right)\|H(x_{k})-\widehat{H}(x_{k})\|^{2}$
$\qquad+\left(\mu_{F}\eta\alpha_{k}-\tfrac{\eta^{2}L^{2}}{2\tau_{k}}-\alpha_{k}\gamma\right)\|x_{k}-\bar{x}_{k}\|^{2}$
$\geq\left(\tfrac{1}{\alpha_{k}}-\tfrac{\tau_{k}}{2}\right)\left(\tfrac{1}{1+\xi}\right)\|H(x_{k})\|^{2}$
$+\left(\mu_{F}\eta\alpha_{k}-\tfrac{\eta^{2}L^{2}}{2\tau_{k}}-\alpha_{k}\gamma-\tfrac{\alpha_{k}^{2}\gamma^{2}}{\xi}\left(\tfrac{1}{\alpha_{k}}-\tfrac{\tau_{k}}{2}\right)\right)\|x_{k}-\bar{x}_{k}\|^{2}.$ (A8)
To make sure $\left(\tfrac{1}{\alpha_{k}}-\tfrac{\tau_{k}}{2}\right)\geq 0$, we choose $\tau_{k}=\tfrac{1.8}{\alpha_{k}}$, then from (C) we can show that
$\|x_{k+1}-\bar{x}_{k}\|^{2}$
$\leq\|x_{k}-\bar{x}_{k}\|^{2}+\|H(x_{k})\|^{2}+\alpha_{k}^{2}\|e_{k}^{\prime}\|^{2}-2\langle x_{k}-\bar{x}_{k},H(x_{k})\rangle$
$\qquad+2\alpha_{k}\|x_{k}-\bar{x}_{k}\|\|e^{\prime}_{k}\|+2\alpha_{k}\|H(x_{k})\|\|e^{\prime}_{k}\|$
$\leq\|x_{k}-\bar{x}_{k}\|^{2}+\|H(x_{k})\|^{2}+2\alpha_{k}^{2}(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2})-2\langle x_{k}-\bar{x}_{k},H(x_{k})\rangle$
$\qquad+\alpha_{k}^{2}\|x_{k}-\bar{x}_{k}\|^{2}+2(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2})+\|H(x_{k})\|^{2}+2\alpha_{k}^{2}(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2})$
$\leq\left(1-2\mu_{F}\eta\alpha_{k}+\tfrac{1}{1.8}\eta^{2}L^{2}\alpha_{k}+\alpha_{k}^{2}+2\alpha_{k}\gamma+\frac{0.2\alpha_{k}\gamma^{2}}{\xi}\right)\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{0.2}{\alpha_{k}(1+\xi)}-2)\|H(x_{k})\|^{2}$
$\qquad+(4\alpha_{k}^{2}+2)(\|e_{k}\|^{2}+\eta^{2}\|\bar{w}_{k,N_{k}}\|^{2}),$
where in the first inequality we used Cauchy-Schwartz inequality, the inequality follows from applying Young’s inequality and (C), and the last one follows from (C).
Next, choosing $\eta=\frac{1.8\mu_{F}}{L^{2}}$, $\xi=9$ and $\gamma\leq 45\left(\sqrt{1+0.04\frac{\mu_{F}^{2}}{L^{2}}}-1\right)$ and taking conditional expectation, one can obtain
$\mathbb{E}[\|x_{k+1}-\bar{x}_{k}\|^{2}\mid\mathcal{F}_{k}]$ $\leq(1+\alpha_{k}^{2})\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{0.02}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\qquad+2(2\alpha_{k}^{2}+1)\big((\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)})^{2}+\tfrac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta^{2}}\nu^{2}}{N_{k}}\big),$
@@
$\mathbb{E}[\|x_{k+1}-\bar{x}_{k+1}\|^{2}\mid\mathcal{F}_{k}]$ $\leq(1+\alpha_{k}^{2})\|x_{k}-\bar{x}_{k}\|^{2}-(\tfrac{0.02}{\alpha_{k}}-2)\|H(x_{k})\|^{2}$
$\qquad+2(2\alpha_{k}^{2}+1)\big((\tfrac{C\rho^{k}}{(k+1)\log^{2}(k+2)})^{2}+\tfrac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\eta^{2}}\nu^{2}}{N_{k}}\big).$
Since $\sum_{k=0}^{\infty}\tfrac{1}{{N_{k}}}<\infty$, by choosing $\alpha_{k}=\tfrac{0.01}{(k+1)^{0.5+\delta}}$ it is easy to show that the requirements of Lemma 3 are satisfied, and we can conclude that (i) $\|x_{k}-\bar{x}_{k}\|^{2}$ is a convergent sequence. Hence, the sequence $\{x_{k}\}_{k}$ is bounded, which guarantees the existence of a convergent subsequence $x_{k_{q}}\to x^{\#}$ as $q\to\infty$, for some $x^{\#}\in X$. (ii) $\sum_{k=0}^{\infty}(\tfrac{0.02}{\alpha_{k}}-2)\|H(x_{k})\|^{2}<\infty$ almost surely. From the definition of $H(x_{k})$, it follows that $\sum_{k=0}^{\infty}2\alpha_{k}(x_{k}-\mathcal{P}_{K(x_{k})}[x_{k}-\eta F(x_{k})])<\infty$. Since $\sum_{k=0}^{\infty}\alpha_{k}=\infty$, we conclude that $\lim_{k\to\infty}x_{k}-\mathcal{P}_{K(x_{k})}[x_{k}-\eta F(x_{k})]\to 0$, which in turn implies that $\lim_{q\to\infty}x_{k_{q}}-\mathcal{P}_{K(x_{k_{q}})}[x_{k_{q}}-\eta F(x_{k_{q}})]\to 0$. Now considering (i) and (ii) we can show that $x^{\#}-\mathcal{P}_{K(x^{\#})}[x^{\#}-\eta F(x^{\#})]=0$, which means that $x^{\#}$ is a solution of SQVI. Therefore, we conclude that the sequence $\{x_{k}\}$ has at least one limit point, and every limit point is a solution of the SQVI almost surely. ∎